OpenStudy (anonymous):

In Triangle ABC, where angle ACB is a right angle. AC = 4 com, AB = 2BC. Find AB. Please help me.. I'm seriously confused D;

OpenStudy (anonymous):

*cm

OpenStudy (anonymous):

Here: use the Pythagoras theorem to obtain : 4^2 + BC^2 = (2BC)^2 or, 16 + BC^2 = 4BC^2 or,3BC^2 = 16 or,BC^2 = 16/3 or,BC = 4/root3 Therefore,AB = 2BC = 8/root3

OpenStudy (anonymous):

I dont understand the part after: or,BC^2 = 16/3.. Why does it become 4/root3 :/

OpenStudy (anonymous):

square root 16 = 4

OpenStudy (anonymous):

OH ok thx a lot :D